### v1

# Safeguard Fine-Tuned LLMs Through Pre- and Post-Tuning Model Merging

Abstract
Fine-tuning large language models (LLMs) for downstream tasks is a widely adopted approach, but it often leads to safety degradation in safety-aligned LLMs. Currently, many solutions address this issue by incorporating additional safety data, which can be impractical in many cases. In this paper, we address the question: How can we improve downstream task performance while preserving safety in LLMs without relying on additional safety data? We propose a simple and effective method that maintains the inherent safety of LLMs while enhancing their downstream task performance: merging the weights of pre- and post-fine-tuned safety-aligned models. Experimental results across various downstream tasks, models, and merging methods demonstrate that this approach effectively mitigates safety degradation while improving downstream task performance, offering a practical solution for adapting safety-aligned LLMs.
Safeguard Fine-Tuned LLMs Through Pre- and Post-Tuning
Model Merging
Hua Farn   Hsuan Su   Shachi H Kumar Saurav Sahay   Shang-Tse Chen   Hung-yi Lee National Taiwan University   Intel Lab alhena.farn@gmail.com
1 Introduction
The rapid advancement and increasing accessibility of Large Language Models (LLMs) necessitate a critical focus on aligning these technologies with human values, cultural norms, and trustworthiness Huang et al. (2023). To address these challenges, researchers and developers have introduced safety techniques such as preference tuning Ouyang et al. (2022); Rafailov et al. (2023); Grattafiori et al. (2024); OpenAI et al. (2024), aimed at preventing LLMs from generating harmful or inappropriate content. Consequently, an increasing number of applications leverage safety-aligned models—referred to as base models in this paper—that are further customized for downstream tasks via Supervised Fine-Tuning (SFT) Chung et al. (2024).
Figure 1: Beyond standard SFT for downstream task adaptation, our proposed method effectively mitigates safety degradation by combining the base and the fine-tuned model while simultaneously enhancing downstream task performance.
However, recent studies Yang et al. (2023); Qi et al. (2024); Zhan et al. (2024) highlight a critical challenge: fine-tuning base models can degrade their safety, even when using benign datasets. To address this issue, mainstream approaches often incorporate additional safety data during fine-tuning Qi et al. (2024); Bianchi et al. (2024). Additionally, some methods that utilize model merging techniques Hazra et al. (2024); Hsu et al. (2024) rely on auxiliary safety or harmful datasets in their complex re-alignment processes. These approaches are often constrained by the scarcity of safety data and can be resource-intensive, requiring datasets with labeled safe examples in specific domains.
Inspired by Wortsman et al. (2022b), we propose a simple yet effective method to improve downstream task performance while mitigating safety degradation. As shown in Figure 1, our approach consists of two steps: first, we fine-tune the base model on the downstream task; second, we merge the base model with the fine-tuned model. We evaluate this approach across various models, downstream tasks, and merging techniques. The experimental results demonstrate that our method enhances downstream task performance while significantly preserving model safety, offering an easy and robust solution for fine-tuning safety-aligned LLMs. Our key contributions are as follows: Our key contributions are as follows:
  • We propose a simple and effective method that improves downstream task performance while reducing the Attack Success Rate (ASR) by up to 30%.
  • We conduct a comprehensive evaluation using three models, four downstream tasks, four merging methods, and two safety benchmarks, demonstrating the robustness of our method in preserving model safety.
2 Related Work
2.1 Jailbreaking LLMs through Fine-Tuning
Despite significant efforts to ensure the safety of Large Language Models (LLMs) Ouyang et al. (2022); Rafailov et al. (2023); Grattafiori et al. (2024); OpenAI et al. (2024), adversarial attacks known as "jailbreaking" remain a major threat Xu et al. (2024); Yi et al. (2024). Additionally, studies Yang et al. (2023); Qi et al. (2024); Zhan et al. (2024) have shown that fine-tuning safety-aligned LLMs can degrade their safety, enabling these models to generate harmful responses.
A common mitigation strategy involves incorporating additional safety data during fine-tuning Qi et al. (2024); Bianchi et al. (2024). Other approaches use complex training or post-training techniques Huang et al. (2024b); Zong et al. (2024); Huang et al. (2024c). However, these methods are often limited by data availability and high computational costs. To overcome these challenges, we propose a simple yet effective method that mitigates safety degradation by merging the base model with the fine-tuned model, without requiring additional safety data or significant computational resources.
2.2 Model Merging
Model merging combines multiple models into a single unified model. A straightforward method averages the weights of different models Wortsman et al. (2022a). Advanced techniques such as SLERP White (2017), DARE Yu et al. (2024), and Model Stock Jang et al. (2024) have been developed to enhance the merging process.
Another approach leverages task vectors Ilharco et al. (2023), enabling flexibility and composability through arithmetic operations for various applications Huang et al. (2024a); Su et al. (2024). Some studies use task vectors to improve LLM safety by fine-tuning additional safe or harmful models to extract a safety vector Bhardwaj et al. (2024); Hazra et al. (2024); Hsu et al. (2024) and determining merging ratios based on safety data Hammoud et al. (2024). In contrast, our method utilizes the inherent safety of the base model, significantly reducing the need for additional safety data.
3 Methodology
To maintain the safety of safety-aligned LLMs after fine-tuning without relying on additional safety data, we propose a two-step approach. First, we fine-tune the base model on a target downstream task. Second, we merge the base model with the fine-tuned model by interpolating their weights. This merging step effectively mitigates safety degradation introduced during fine-tuning by leveraging the original safety-aligned LLMs.
Step 1: Fine-Tuning the Large Language Model
We fine-tune the base model on downstream tasks. For each task $t$, given an instruction $x^{t}$ and the corresponding response $y^{t}$, we optimize the model using the following loss function:
$\mathcal{L}_{FT}=-\log f_{\theta}(y^{t}|x^{t})$
where $f$ represents the language model parameterized by $\theta$.
Step 2: Merging the fine-tuned LLM with the base model
After fine-tuning, we merge the base model parameters ($\theta_{base}$) with those of the fine-tuned model ($\theta_{t}$) using the interpolation formula:
$\theta_{merged}=(1-\lambda)\theta_{base}+\lambda\theta_{t}$
Here, $\theta_{merged}$ represents the parameters of the merged model, and $\lambda\in[0,1]$ determines the contribution of the fine-tuned model. The ratios are normalized to ensure they sum to 1. In our experiments, we explore various merging techniques based on Equation 3, including methods that adjust model parameters before merging and strategies to identify the optimal $\lambda$.
4 Experimental Setups
In this section, we describe the experimental setup to address the following research questions: Q1: Can merging fine-tuned models with their base models prevent safety degradation after fine-tuning on different downstream tasks? Q2: How do different merging methods perform? Q3: What is the impact of $\lambda$ on downstream task performance and safety? Q4: Does model merging work on multiple models?
4.1 Downstream Tasks
We evaluate our proposed method on four downstream tasks: reasoning, medical assistance, code generation, and tool usage proficiency.
Reasoning is improved with Chain-of-Thought data from the Flan V2 dataset Longpre et al. (2023) and is evaluated on the Big Bench Hard (BBH) dataset Suzgun et al. (2023). Medical assistance uses patient-doctor interactions from the ChatDoctor dataset Li et al. (2023). Code generation uses the MagiCoder dataset Wei et al. (2024) and is assessed via the HumanEval dataset Chen et al. (2021). Tool usage proficiency uses the OpenFunctions dataset Patil et al. (2023) for enhancing API call generation. For medical assistance and tool usage proficiency, response similarity to reference answers is measured using BERTScore [^1] Embedding extraction from the 40th layer of microsoft/deberta-xlarge-mnli. Zhang* et al. (2020). See Appendix A for more details on the downstream tasks.
4.2 Safety Evaluation
Safety is assessed using harmful instructions from the AdvBench dataset Chen et al. (2022) and the HEx-PHI dataset Qi et al. (2024). AdvBench contains 520 harmful instructions and HEx-PHI comprises 330 harmful instructions. We employ the WildGuard safety classification model Han et al. (2024), which demonstrates performance comparable to GPT-4 OpenAI et al. (2024). The Attack Success Rate (ASR) is used as the metric for both datasets. Further details on safety evaluation are provided in Appendix B.
4.3 Large Language Models
Our experiments involve two LLM families: Llama-3 Grattafiori et al. (2024) and Gemma Team et al. (2024). We use their instruct-tuned versions, which are aligned with human preferences.
Within the Llama-3 family, we evaluate Llama-3-8B-Instruct and Llama-3.1-8B-Instruct. For the Gemma family, we use the Gemma-2B-it model. Each model is fine-tuned with three different seeds using LoRA Hu et al. (2022) for each downstream task. Additional details on training and inference setups are provided in Appendix C.
4.4 Merging Methods
We evaluate three merging methods: Linear Merging, where models are directly combined by interpolating their weights following Equation 3, Model Stock Jang et al. (2024), SLERP White (2017), and DARE Yu et al. (2024). For Linear Merging, SLERP, and DARE, each fine-tuned model is merged with the base model using $\lambda$ selected based on validation set performance on the downstream tasks. Specifically, in the Model Stock approach, three models fine-tuned with different seeds are merged with the base model. More details about the methods and implementations are provided in Appendix D.
5 Results and Discussion
In this section, we present the results addressing our research questions. From Q1 through Q4, we focus primarily on the Llama-3-8B-Instruct model, other models are discussed in the context of Q5.
Q1: Can merging fine-tuned models with their base models prevent safety degradation after fine-tuning on different downstream tasks?
[…]
Table 1: Results across four tasks. The table reports the average scores of three models, which are fine-tuned with different seeds, for FT and Linear Merging. Performance indicates task-specific effectiveness measured by the respective metrics.
The results for Llama-3-8B-Instruct are presented in Table 1. While the fine-tuned model (FT) demonstrates improved performance on the downstream task, it significantly compromises safety. However, merging the fine-tuned model with the base model substantially enhances safety across all downstream tasks, as highlighted in the green cells in Table 1 and subsequent tables. Although the merged model does not fully match the base model in safety, it substantially reduces the ASR on both the AdvBench and HEx-PHI safety datasets. Notably, on the medical assistance task, the ASR is reduced by over 30%, bringing safety closer to the base model’s level. This result underscores the base model’s stronger inherent safety features. Additionally, Linear Merging improves downstream task performance across all downstream tasks, consistent with the findings of Wortsman et al. (2022b). Given the pronounced effect on the medical assistance task, subsequent experiments for questions Q2 through Q4 primarily focus on this task.
Q2: How do different merging methods perform?
[…]
Table 2: Results of various merging methods on the medical assistance task. The results show that different merging methods are beneficial for the safety of the fine-tuned model while also enhancing task performance.
The performance of alternative merging methods is shown in Table 2. Both SLERP and DARE effectively reduce ASR on AdvBench and HEx-PHI. The results emphasize the strong performance of Linear Merging, demonstrating that simpler approaches can effectively mitigate safety degradation. This finding suggests that Linear Merging can be a viable and computationally efficient alternative for practical applications.
Q3: What is the impact of $\lambda$ on downstream task performance and safety?
Figure 2: Impact of $\lambda$ on downstream task performance and safety. These plots illustrate how varying $\lambda$ affects the performance of the Medical Assistance task (left y-axis) and the ASR for the HEx-PHI dataset (right y-axis) across three merging methods. Results for AdvBench are shown in Figure 3 in the Appendix.
[…]
Table 3: Results of Llama-3.1-8B-Instruct and Gemma-2B-it The results show that our method can generalize to more models, restoring safety while maintaining task performance.
Figure 2 illustrates the impact of $\lambda$ on downstream task performance and model safety. As $\lambda$ increases, downstream task performance improves, but the ASR also rises, indicating a trade-off between performance and safety. However, the ASR remains lower than that of the SFT model. The optimal $\lambda$ is observed around $0.5\sim 0.6$, suggesting that more evenly combining the weights of the two models yields better downstream task performance while maintaining safety. Additionally, the rate of ASR increase for Linear Merging, as shown in Figures 2 and 3, is slower than that of SLERP and DARE, further demonstrating that Linear Merging is a more practical method in real-world applications.
Q4: Does model merging work on multiple models?
To evaluate the generalizability of model merging, we test Llama-3.1-8B-Instruct and Gemma-2B-it, with results shown in the Table 3. For these models, safety degradation on both benchmarks after fine-tuning is relatively mild. Nevertheless, model merging effectively restores safety without significantly compromising downstream task performance, demonstrating the method’s applicability across different LLMs.
6 Conclusion
We propose a two-step method to address safety degradation in aligned LLMs by merging the weights of pre- and post-fine-tuned models. This approach allows aligned LLMs to acquire new downstream task capabilities while preserving their original safety features, without the need for additional safety data. Experiments demonstrate its effectiveness across diverse downstream tasks, models, and merging techniques, providing a simple yet robust solution for adapting aligned LLMs.
7 Limitations
Task and Model Selection
In our experiments, we evaluate only on benign data from the reasoning, medical assistance, code generation, and tool-using proficiency tasks, leaving out other areas such as law, finance, and others. While Section 5 examines the efficacy of our methods on these four downstream tasks, the performance of aligned models fine-tuned on other domains, languages, or datasets contaminated with harmful examples remains an area for future investigation. Furthermore, our methods are tested only on models with sizes of 2B and 8B from two model families. The efficacy of the approach on larger models and different model families is uncertain and warrants further exploration.
Safety Classifier for Safety Evaluation
Due to the higher cost of methods like LLM-as-Judge Chiang and Lee (2023); Liu et al. (2023), which involve using LLMs such as GPT-4 OpenAI et al. (2024) to evaluate the safety of model responses, we utilize an alternative approach by employing a classifier, WildGuard Han et al. (2024), to classify model responses as safe or unsafe. While this reduces costs, WildGuard has limitations: it struggles with more complex instructions, may produce false positives or false negatives, and provides less detailed evaluations. Consequently, we are unable to analyze which types of harmful instructions the models are more vulnerable to or which types are most effectively defended against after applying our method. More detailed safety analysis is left for future work.
8 Ethics Statement
While our method effectively addresses safety degradation in aligned LLMs without requiring additional safety data, our approach relies on merging pre- and post-fine-tuned models to preserve safety, which may inadvertently inherit any latent biases or unsafe behaviors that are still presented in the base model. Further in vestigation is needed to explore the impact of these inherited biases in the base model.
Appendix A Domain-Specific Tasks Detail
Reasoning
We randomly select 10,000 zero-shot chain-of-thought instructions from the Flan V2 dataset then split them into training set and validation set with ratio $9:1$. Performance is assessed using the BBH dataset, with results reported as the average 3-shot accuracy across all BBH tasks. We use lm-evaluation-harness Gao et al. (2024) as our code base.
Medical Assistance
We randomly select 10,000 real patient-doctor conversations from the ChatDoctor dataset Li et al. (2023) then split them into training set and validation set with ratio $9:1$. Model performance is evaluated on 1,000 unseen patient queries using BERTScore to calculat similarity of reference responses and models’ responses, we report the F1 score in our results.
Code Generation
We select 10,000 samples from the MagiCoder dataset Wei et al. (2024) to improve code generation capabilities. Specifically, we uniformly sampled from each coding languages. We report Pass@10 in our experiment results.
Tool Using Proficiency
Due the the smaller dataset size of OpenFunctions Patil et al. (2023), we split its full training set into two subsets with ratio $9:1$ for training and validating to enhance API call generation capabilities. The model is evaluated on the full OpenFunctions test set, with performance measured using BERTScore to calculate similarity of reference responses and models’ responses, and we report the F1 score in our results.
Appendix B Safety Dataset and Classifier
B.1 Safety Dataset
AdvBench
Chen et al. (2022) is a set of 520 harmful behaviors formulated as instructions. The goal of these instructions is to identify a single attack string that causes the model to generate responses compliant with the instruction across as many harmful behaviors as possible.
HEx-PHI
Qi et al. (2024) is a dataset that contains 330 harmful instructions for evaluating the harmfulness of LLMs. It comprises 30 examples for each of 11 prohibited categories, such as Illegal Activity, Child Abuse Content, Hate/Harass/Violence, and so on.
B.2 Safety Classifier
WildGuard Han et al. (2024) is an open, one-stop moderation model that achieves three goals: 1) Detection of harm in user prompts, 2) Detection of harm in LLM responses, and 3) Refusal evaluation of LLM responses. Comprehensive evaluations on WildGuardTest Han et al. (2024) and ten existing public benchmarks that are shown in Han et al. (2024) demonstrate that WildGuard outperforms the strongest existing open-source baselines in F1 scores and matches GPT-4 across various tasks.
When evaluating the responses generated by LLMs to harmful instructions, we apply the recommended settings, such as the instruction format and decoding strategy of WildGuard [^2] https://huggingface.co/allenai/wildguard. After evaluation, WildGuard outputs the response in the following format:
Harmful request: yes/no
Response refusal: yes/no
Harmful response: yes/no
We then parse the result of the harmful response. If the result cannot be parsed, we count it as a miss. Our final ASR is calculated as follows:
$ASR=\frac{R_{yes}}{R_{total}-R_{miss}}$
where $R_{yes}$ is the number of harmful responses classified as "yes", $R_{total}$ is the total number of responses, and $R_{miss}$ is the number of responses that failed to be parsed. In our experiments, $R_{miss}$ is usually less than 5 for all tested models across both safety datasets.
Appendix C Experimental Detail
C.1 Prompt Template
For base models, we directly apply their own prompt templates during the training and inference phases. For fine-tuned models, we apply the prompt templates of their respective base models.
For the Llama-3 family, we use the following prompt template with a system prompt: You are a helpful assistant. for the tasks of reasoning, code generation, and tool usage proficiency:
<|begin_of_text|><|start_header_id|>system<|end_header_id|>
You are a helpful assistant.<|eot_id|>
<|start_header_id|>user<|end_header_id|>
{Instruction}<|eot_id|>
<|start_header_id|>assistant<|end_header_id|>
{Response}
For the medical assistance task, we use another prompt provided in the ChatDoctor dataset Li et al. (2023) as the system prompt. Hence, the prompt is as follows:
<|begin_of_text|><|start_header_id|>system<|end_header_id|>
If you are a doctor, please answer the medical
questions based on the patient’s description.<|eot_id|>
<|start_header_id|>user<|end_header_id|>
{Instruction}<|eot_id|>
<|start_header_id|>assistant<|end_header_id|>
{Response}
The prompt for Gemma-2B-it for the tasks of reasoning, code generation, and tool usage proficiency is shown below:
<bos><start_of_turn>user
You are a helpful assistant.{Instruction}<end_of_turn>
<start_of_turn>model
{Response}
The prompt for the medical assistance task is as follows:
<bos><start_of_turn>user
If you are a doctor, please answer the medical
questions based on the patient’s description.
{Instruction}<end_of_turn>
<start_of_turn>model
{Response}
C.2 Fine-tuning
For all tasks, we fine-tune three model instances using different random seeds: 42, 1024, and 48763. We employ LoRA with $r=8$ and $\alpha=16$ for all linear layers, utilizing the AdamW optimizer with a learning rate of $1\times 10^{-4}$ and a cosine learning rate scheduler. We use a batch size of 8 and train for 3 epochs. All models are trained on either an RTX A6000 GPU or an RTX 6000 Ada Generation GPU using LLaMA-Factory Zheng et al. (2024) as the codebase.
Although we initially fine-tuned each task for 3 epochs, we observed stronger model performance at an earlier stage. Consequently, unless explicitly stated otherwise, we report model training after 500 steps for reasoning, medical assistance, and code generation, and after 200 steps for tool usage proficiency due to the smaller size of the OpenFunctions training set.
C.3 Inference
We use greedy decoding to ensure result consistency, except for the HumanEval benchmark. For HumanEval, we apply sampling-based decoding with a temperature of 0.6, top_p of 0.9, top_k of 50, and a repetition penalty of 1.2. To accelerate the inference process, we utilize the VLLM engine for model inference.
Appendix D Model Merging
D.1 Merging Methods
Linear Merging
Linear Merging involves directly combining the weights of the base model and the fine-tuned model by interpolating their parameters. Specifically, the weights of the merged model are calculated as a weighted average of the base and fine-tuned models’ weights, following Equation 3. This method is straightforward and computationally efficient, making it a popular choice for basic model integration.
SLERP
Spherical Linear Interpolation (SLERP) White (2017) is an advanced merging technique that interpolates between model weights on a hypersphere, ensuring a smoother and more natural transition between the models. Unlike Linear Merging, SLERP accounts for the angular relationship between weight vectors, which aim to better preserve the base model’s features while effectively integrating the fine-tuned model’s task-specific enhancements.
DARE
Drop and Rescale (DARE) Yu et al. (2024) is a method used to prepare models for merging techniques such as Linear Merging. It operates by randomly dropping parameters according to a specified drop rate and rescaling the remaining parameters. This process helps reduce the number of redundant and potentially interfering parameters among multiple models. In our experiments, we apply DARE only to the fine-tuned models and then use Equation 3 to merge them with the base model.
Model Stock
Model Stock Jang et al. (2024) utilizes the geometric properties of the weights of fine-tuned models. It determines the optimal merging ratio for each model by minimizing the distance between the merged weights and the center of the weights of the fine-tuned models and the base model. However, since this method requires at least two fine-tuned models and a base model, we choose to merge three models with the base model in our experiments. Therefore, when implementing Model Stock, we first uniformly average the weights of the three fine-tuned models, denoted as
$\theta_{m}=\frac{\theta_{42}+\theta_{1024}+\theta_{48763}}{3},$
where $\theta_{42},\theta_{1024},\theta_{48763}$ represent the model trained with different seed, then we merge this average with the base model using the optimal ratio $\lambda_{optimal}$ determined by Model Stock. The weight of the final merged model is expressed as:
$\theta_{merged}=(1-\lambda_{optimal})\theta_{base}+\lambda_{optimal}\theta_{m}.$
D.2 Model Merging Implementation
For Linear Merging, we determine the interpolation weight $\lambda$ by testing values in the set $\{0.2,0.4,0.6,0.8\}$. For SLERP and DARE, we test the interpolation weight $\lambda$ in the same set, $\{0.2,0.4,0.6,0.8\}$, but do not experiment with the threshold for the dot product in the SLERP algorithm. Regarding Model Stock, since it can automatically approximate the interpolation weight based on the geometric nature of the training trajectory, no hyperparameters need to be specified. We use MergeKit Goddard et al. (2024) as our codebase.
Appendix E Impact of $\lambda$ for Safety
We also illustrate the impact of $\lambda$ on AdvBench in Figure 3. Similar to the results for HEx-PHI, the ASR increases as $\lambda$ increases, but it remains better than the results of the SFT model.
Figure 3: Impact of $\lambda$ on Task Performance and Safety. This figure illustrates how varying $\lambda$ affects the ASR for the AdvBench dataset across different merging methods.